#1 The deli charges $3.80 for a breakfast of 3 eggs & 4 bacon strips. The charge is $2.75 for 2 eggs & 3 bacon strips. Find the cost of each egg and the cost of each bacon strip?
(Please help me I don't even no where to begin with these I freeze up.)
#2 If 3/4 is added to three times a number, the result is 1/2 subtracted from twice the number. Find the number?
#3 Carol spent $32.25 to take her daughter's birthday party guest to the movies. Adult tickets cost $5.75 and children tickets cost $3.00. If 8 people were at the party.how many of each ticket were bought?
#4 A chemist working on a new project needs 12 liters of a 50% acid solution. The stockroom only has 40% and 70% solutions. How much of each solution should be mixed together to form 12 liters of a 50% solution?
# 5 Solve: 2x-(3x-2)+5=17

Answer by mananth(16949)

1 The deli charges $3.80 for a breakfast of 3 eggs & 4 bacon strips. The charge is $2.75 for 2 eggs & 3 bacon strips. Find the cost of each egg and the cost of each bacon strip?
(Please help me I don't even no where to begin with these I freeze up.)
let egg = $x
bacon = $y
3x+4y=3.80...............1
2x+3y=2.75...............2
3 x + 4 y = 3.8 .............1
2 x + 3 y = 2.75 .............2

multiply (1)by -2
Multiply (2) by 3
-6x-2y =-7.6
6x+3y=8.25

Add the two equations

y=0.65

Plug value of y in
3 x + 4 y = 3.8
3 x + 2.6 = 3.8
3 x + = 1.2
/ 3
x= 0.4
eggs 40 cents, bacon 65 cents
#2 If 3/4 is added to three times a number, the result is 1/2 subtracted from twice the number. Find the number?
Let x be the number
3x+3/4 =2x - 1/2
3x-2x=-1/2 -3/4
x=-5/4

#3 Carol spent $32.25 to take her daughter's birthday party guest to the movies. Adult tickets cost $5.75 and children tickets cost $3.00. If 8 people were at the party.how many of each ticket were bought?
adults x
children y
x+y =8..............1
5.75x+3y= 32.25.....2
1 x + 1 y = 8 .............1
5.75 x + 3 y = 32.25 .............2

multiply (1)by -3
Multiply (2) by 1
-3x-3y =-24
5.75x+3y =32.25

Add the two equations

2.75x=8.25

/2.75

x=3

plug value of x in (1)

1x+1y=8
3+1y=8
y=8-3
y=5

#4 A chemist working on a new project needs 12 liters of a 50% acid solution. The stockroom only has 40% and 70% solutions. How much of each solution should be mixed together to form 12 liters of a 50% solution?
percent ---------------- quantity
40 ---------------- x
70 ---------------- 12-x
50.00% ---------------- 12
40x+70(12-x)=50* 12
40x+840 -70x=600
40x-70x =600-840
-30x=-240
/-30
x=8liters 40%
# 5 Solve: 2x-(3x-2)+5=17
2x-3x+2+5=17
-x+7=17
-x=10
x=-10